OpenStudy (nuccioreggie):

please help me Use the rules for integer addition to simplify the expression. Positive twelve plus negative eleven equals what value? negative twenty-three negative one positive one positive twenty-three

OpenStudy (nuccioreggie):

@jim_thompson5910

OpenStudy (math&ing001):

(+12) + (-11) = 12-11

OpenStudy (mathstudent55):

These are the rules of integer addition. 1. To add two positive numbers, just add them as you've been doing since elementary school. 2. To add two negative numbers, remove the negative signs. Add the numbers as positive numbers. Then make the result negative. 3. To add a positive number and a negative number, take the absolute value of both numbers. Then subtract the smaller absolute value from the larger absolute value. The answer has the same sign as the sign of the number with the larger absolute value.

OpenStudy (mathstudent55):

Now let's look at your question. +12 + (-11) We are adding a positive number and a negative number. This is case 3 above. a. Take the absolute values: 12 and 11 b. Subtract the smaller abs value from the larger abs value: 12 - 11 = 1 c. The answer has the same sign as the number with the larger absolute value. 12 is greater than 11. 12 is positive, so our answer is positive. The answer is 1. +12 + (-11) = 1

OpenStudy (nuccioreggie):

oooooo so the correct answer is positive 1 ??

OpenStudy (mathstudent55):

yes